OpenStudy (anonymous):

A company installs 5000 light bulbs, each with an average life of 500 hours, standard deviation of 100 hours, and distribution approximated by a normal curve. Find the percentage of bulbs that can be expected to last the period of time. Less than 690 hours A. 97.06% B. 47.14% C. 97.1% D. 97.2%

OpenStudy (anonymous):

how do i put this in a math question

OpenStudy (anonymous):

Hello, 1) Since the mean is 500, 1/2 of them would last less than 500 hr. 2) 540 is .4 SD from the mean n = (540 -500) /100 = 0.4 so we look up z of 0.4 = 0.6554 780 is 2.8 SD from the mean m = (780-500)/100 = 2.8 looking up z of 2.8 gives us 0.9974 So we want the percent in between these hence: 0.9974 - 0.6554 = 0.3420 as a percent 34.20% Hope This Helps! http://answers.yahoo.com/question/index?qid=20110506120154AA68QFZ